# Hausman’s Consistency Test and an Internal Bias Diagnostic: A Short Note Regarding Linear Mixed Models

### v1

@@ Abstract.
The Hausman specification test detects inconsistency in mixed model estimators of fixed effects by comparing the original model with an alternative specification in which random effects are treated as fixed. This note illustrates a bias diagnostic from the statistical literature as a complement to the Hausman test. The diagnostic can provide internal estimates of parameter-specific bias in the mixed model estimators without requiring the second, fixed-effects-only model to be fit. We apply the diagnostic to a panel data analysis as well as to a Value-Added Model (VAM) for teacher evaluation.
@@
linear mixed model, misspecification, permutation test, stochastic model matrix
@@ 1. Introduction
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},{\mathbf{R}})$. For the resulting estimator of $\boldsymbol{\beta}$ to be consistent, it is necessary that the random effects $\boldsymbol{\eta}$ are uncorrelated with the regressors in ${\mathbf{X}}$ (Hausman & Taylor, 1981a). The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-$\boldsymbol{\eta}$ effects (RE) and fixed-$\boldsymbol{\eta}$ effects (FE) estimators, the latter requiring a re-fit of the model that treats $\boldsymbol{\eta}$ as an additional vector of fixed effects. Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them. Under the alternative, only the FE estimator is consistent. A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
@@
The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity. By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size. While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias. To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification.
@@
We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when the random-effects design matrix ${\mathbf{Z}}$ may be correlated with the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based p-value, all without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability. For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
@@
In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test. While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between ${\mathbf{Z}}$ and $\boldsymbol{\eta}$. Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
@@ 2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting ${\mathbf{M}}^{-}$ represent a generalized inverse of ${\mathbf{M}}$, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})^{\prime}[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})]^{-}(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})$
is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})$ (Hausman & Taylor, 1981a). A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
@@ 2.2. Bias Diagnostic
The RE estimator is
$\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. Under the assumption that ${\mathbf{X}}$ is fixed and ${\mathbf{Z}}$ is stochastic, Karl & Zimmerman (2021) show
$\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$
where
$\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}$
@@
and that this bias is zero if any of the following three conditions are true: (1) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (2) if $\boldsymbol{\eta}$ and ${\mathbf{Z}}$ are independent, or (3) if $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator). The third condition indicates that the FE estimator remains unbiased in the presence of model misspecification that biases the RE estimator. Thus the internal bias estimate often approximates the difference between the RE and FE estimators.
The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal bias estimate for the RE estimator. A vector permutation function $\boldsymbol{\pi}(\cdot)$ is applied repeatedly to calculate draws $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}})$ from a reference distribution to assess if the observed bias is unusually large under the null hypothesis that $\boldsymbol{\eta}$ is independent of all other model components. The bias p-value is calculated as the proportion of reference values that are larger in absolute value than that of the internal plug-in bias estimate. This test has been implemented in the R package mixedbiastest (Karl, 2024): the package accepts mixed model objects from the popular lme4 package (Bates et al., 2015) and produces bias diagnostics for each parameter of $\boldsymbol{\beta}$ as well as custom contrasts. mixedbiastest currently only accepts lme4 models with diagonal random effects covariance matrices $({\mathbf{G}})$, but this matches the structure of most panel data applications and could be generalized as described by Karl & Zimmerman (2021).
@@ 3. Gasoline Consumption Data Application
We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test.
@@
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test
@@
Figure 1.  Graphical Summary of Bias Diagnostics for Gasoline Data. The histograms show the reference distributions, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}}),$ and the dashed red lines show the internal bias estimates, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$.
@@
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test indicated inconsistency of the RE estimator, and bias diagnostic now suggests that lrpmg is significantly biased downward. Note the similarity of the internal bias estimates from the RE model with the RE-FE differences in Table 1.
@@ 4. Value-added Model (VAM) Results with GPvam
Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014). There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix (${\mathbf{R}}$). Classroom-heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$). A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix ${\mathbf{G}}$ with the number of unique elements equal to the number of grades in the study. The random effects model matrix ${\mathbf{Z}}$ tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple member” structure for ${\mathbf{Z}}$ (Browne et al., 2001). The nonrandom assignment of students to classrooms creates the potential for biasing (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021). Calculating the Hausman test statistic to test consistency for this example would require customized programming to fit a FE model that treats the teacher effects as fixed while still utilizing the multiple membership structure of ${\mathbf{Z}}$ and the block diagonal structure of the error covariance matrix, $\mathbf{R}$. By contrast, the bias test only requires the estimates and model matrices from the fitted mixed model.
The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic. Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the ${\mathbf{Z}}$, ${\mathbf{G}}$, and $\mathbf{R}$ matrices can be constructed. This is implicit in the function call, and the formula argument to GPvam only constructs the fixed effects structure ${\mathbf{X}}$ and $\boldsymbol{\beta}$. We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6. This is the same data set used by Karl et al. (2013a).
@@
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased. The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a p-value of 0 (none of the one million permutations exceeded the test statistic). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis.

### v2

# Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models

@@ Abstract.
The Hausman specification test detects inconsistency in random-effects estimators by comparing them with alternative fixed-effects estimators. This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications. The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model. We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
@@
linear mixed model, Hausman test, correlated random effects, permutation test
@@ 1. Introduction
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},\mathbf{R})$. Here $\mathbf{y}$} is an $n\times 1$ response vector, $\mathbf{X}$} is an $n\times p$ fixed-effects design matrix with associated coefficient vector $\boldsymbol{\beta}$, and $\mathbf{Z}$} is an $n\times m$ random-effects design matrix with random-effects vector $\boldsymbol{\eta}$. The matrices $\mathbf{G}$} and $\mathbf{R}$ denote the $m\times m$ random-effects covariance matrix and the $n\times n$ error covariance matrix, respectively.
@@
In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors (the usual exogeneity condition; see, e.g., Hausman & Taylor, 1981a). The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-effects (RE) and fixed-effects (FE) estimators. In its original formulation, this involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests based on a single augmented RE fit; see Section 2. Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them. Under the alternative, only the FE estimator is consistent. A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
@@
The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity. By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size. While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias. To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification. In what follows, we deliberately keep these roles separate: the Hausman test is used as a classical large-sample specification test for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to quantify the magnitude and direction of parameter-specific bias in the same RE fit.
@@
We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} may be stochastically dependent on the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value, all without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability. For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
@@
In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test. While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between $\mathbf{Z}$} and $\boldsymbol{\eta}$. Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model. Our aim is not to propose a new alternative to the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing linear mixed model bias diagnostic can supplement them by providing coefficient- or contrast-specific information obtained from a single fitted mixed model.
@@ 2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting $\mathbf{M}$$^{-}$ represent a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a). A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent. In the canonical panel-data setting, this statistic is derived under the assumption of homoskedastic, serially uncorrelated errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$. In the linear mixed model formulation we write the model with a general error covariance matrix $\mathbf{R}$; the gasoline example below maintains the standard homoskedastic assumption, whereas the value-added model in Section 4 uses a block-diagonal $\mathbf{R}$.
@@ 2.2. Bias Diagnostic
The random-effects (RE) estimator of $\boldsymbol{\beta}$ is
(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic. Let ${\mathbf{V}}={\mathbf{Z}}{\mathbf{G}}{\mathbf{Z}}^{\prime}+\mathbf{R}$ denote the true marginal covariance matrix of $\mathbf{y}$}, and define
$\boldsymbol{\nu_{k}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}{\mathbf{V}}^{-1}{\mathbf{Z}}.$
Karl & Zimmerman (2021) show that for any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$),
(3) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\boldsymbol{\nu_{k}}^{\prime}\boldsymbol{\eta}].$
Karl & Zimmerman (2021) propose a plug-in approximation based on the fitted mixed model. Writing
(4) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}},$
@@
and denoting the empirical best linear unbiased predictors (EBLUPs) of the random effects by $\hat{\boldsymbol{\eta}}$, the plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal estimate of the bias of $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$. The exact bias is zero if any of the following three conditions hold: (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, or (iii) $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator). The third condition implies that the FE estimator remains unbiased in settings where the RE estimator is biased, so that the internal bias estimate often approximates the difference between the RE and FE estimators (Karl & Zimmerman, 2021).
[…]
2.3. Relation to correlated random effects and Mundlak–Wooldridge tests
In the panel-data literature, a common way to test the random-effects specification is through correlated random effects (CRE) or “Mundlak” type augmentations of the RE model. Suppose we have the panel-data setup
$y_{it}=x_{it}^{\prime}\beta+c_{i}+\varepsilon_{it},$
with individual effect $c_{i}$ and idiosyncratic error $\varepsilon_{it}$. Under the usual random-effects assumptions, including strict exogeneity and $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ (often summarized as $\operatorname{Cov}(x_{it},c_{i})=0$), the random-effects estimator is consistent. Mundlak (1978) models potential correlation between $x_{it}$ and $c_{i}$ by including the group means $\bar{x}_{i}$ in the regression. Wooldridge (2010) shows that a simple Hausman-type alternative is to estimate the augmented random-effects model
$y_{it}=x_{it}^{\prime}\beta+\bar{x}_{i}^{\prime}\gamma+c_{i}+\varepsilon_{it},$
and apply a (robust) Wald test of $H_{0}:\gamma=0$. This CRE test uses only an augmented RE fit but is asymptotically equivalent to the traditional FE–RE Hausman test when the idiosyncratic errors are homoskedastic and serially uncorrelated (Hausman & Taylor, 1981a). Robust implementations based on covariance estimators that are consistent under general forms of heteroskedasticity and serial correlation are widely used in applied work; see, e.g., Wooldridge (2010).
In terms of the linear mixed model (1), stacking the panel outcomes $\{y_{it}\}$ into $\mathbf{y}$} and the regressors $\{x_{it}\}$ into the rows of $\mathbf{X}$} yields ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon}$, where $\boldsymbol{\eta}$ collects the individual effects $c_{i}$ and $\mathbf{Z}$} is the $n\times N$ incidence matrix (with $N$ the number of individuals) that has a one in the row corresponding to observation $(i,t)$ and column $i$ when that observation belongs to individual $i$. In this formulation the assignment structure encoded by $\mathbf{Z}$} may itself be stochastic and potentially correlated with the random effects $\boldsymbol{\eta}$, which is precisely the type of dependence considered by Karl & Zimmerman (2021). In the conventional RE panel-data model, the exogeneity condition $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ restricts the dependence between the random effects $c_{i}$ and the regressors $x_{it}$. By contrast, the zero-bias conditions listed in the previous subsection include a requirement that rules out dependence between the random effects and the random-effects design matrix $\mathbf{Z}$}. In simple random-intercept panels $\mathbf{Z}$} is just an incidence matrix and is often implicitly assumed to be fixed without further consideration.
Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that arises when the random-effects design matrix $\mathbf{Z}$} and the random-effects vector $\boldsymbol{\eta}$ are dependent, with $\mathbf{X}$} treated as fixed. The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ and its permutation-based $p$-value can therefore be viewed as a parameter-specific analogue of a Hausman-type specification check, obtained from a single fitted mixed model. In applications where both $\mathbf{X}$} and $\mathbf{Z}$} may reasonably be regarded as stochastic, the diagnostic still targets the component of bias that operates through the joint behavior of $\mathbf{Z}$} and $\boldsymbol{\eta}$; it does not attempt to separate any additional bias that might arise from stochastic variation in $\mathbf{X}$} alone. In this sense, the Karl & Zimmerman (2021) diagnostic focuses on a narrower class of misspecification than the classical Hausman or Mundlak–Wooldridge tests, and is best viewed as a complementary, parameter-specific tool rather than a replacement for those global specification checks.
An advantage of this approach is that it can be implemented directly from a single fitted mixed model, without requiring the analyst to construct a full FE or CRE analogue in settings that depart from the standard individual-by-time panel, such as models with multiple-membership random effects, nontrivial error-covariance structures $\mathbf{R}$, or a very large number of random-effect levels. In such applications, FE or CRE-style specifications are often still possible in principle, but may require additional data construction and modeling work. The value-added model in Section 4 provides one such example. When conventional FE or CRE estimators are readily available, we recommend using the classical Hausman or Mundlak–Wooldridge tests as global checks of the RE specification, and then using the bias diagnostic to identify which coefficients or contrasts exhibit the largest finite-sample bias and in which direction.
@@ 3. Gasoline Consumption Data Application
We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test. All results in this section can be reproduced using Listings 1 and 2 together with the Gasoline data supplied by the plm package.
@@
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test.
@@
[…]
@@
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test indicates global inconsistency of the RE estimators, and the bias diagnostic suggests that the RE estimate of lrpmg is significantly biased downward. Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
@@ 4. Value-added Model (VAM) Results with GPvam
Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014). There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix ($\mathbf{R}$). Classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$). A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with the number of unique elements equal to the number of grades in the study. The random-effects model matrix $\mathbf{Z}$} tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple-membership” structure for $\mathbf{Z}$} (Browne et al., 2001). The nonrandom assignment of students to classrooms creates the potential for bias (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021), through dependence between the random-effects design matrix $\mathbf{Z}$} and the teacher effects $\boldsymbol{\eta}$.
These features make the CP VAM a natural setting in which to use a parameter-specific bias diagnostic. The random-effects structure involves thousands of teacher effects, the student-level error covariance matrix $\mathbf{R}$ is block-diagonal rather than homoskedastic, and the multiple-membership design encoded by $\mathbf{Z}$} is more complex than a simple random-intercept panel. The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic. Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices can be constructed. This is implicit in the function call, and the formula argument to GPvam only constructs the fixed-effects design matrix $\mathbf{X}$} and parameter vector $\boldsymbol{\beta}$. We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6. This is the same data set used by Karl et al. (2013a). In this value-added model application, we use the GPvam implementation of the bias diagnostic, which accommodates the block-diagonal error covariance matrix and multiple-membership random effects structure that fall outside the scope of the current mixedbiastest implementation.
@@
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased. The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a $p$-value of 0 (none of the one million permutations exceeded the test statistic). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis. In this example we do not propose a change in estimator based solely on the diagnostic. Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.